Nonnuclear submarines use batteries for power when submerged. (a) Find the magnetic field 50.0 cm from a straight wire carrying 1200 A from the batteries to the drive mechanism of a submarine. (b) What is the field if the wires to and from the drive mechanism are side by side? (c) Discuss the effects this could have for a compass on the submarine that is not shielded.
Answer #1

B = magnetic field

i = current = 1200 A

r = distance from the wire = 50 cm = 0.50 m

a)

magnetic field is given as

B = (\mu/4\pi) (2i/r) = (10-7) (2 x 1200/0.5) = 0.48 x 10-3 T